OpenStudy (anonymous):

What are the solutions to 3|x + 2| = 18?

OpenStudy (anonymous):

-8 or 4

OpenStudy (anonymous):

how did u do it

OpenStudy (anonymous):

this is how we do it , if x+2 is positif , so it means that 3x+6=18 ,then 3x=18-6, x=12/3=4 ,if x+2 negative , so -3x-6=18, -3x=24, x=-24 got it ?

OpenStudy (anonymous):

Do you know the definition of absolute value? An absolute value can we translated to a piece-wise form like this:\[\bf \left| x \right|=\left[\begin{matrix}x \ \ \ for \ x \ge 0 \\ \bf -x \ \ \ for \ x<0 \end{matrix}\right]\]

OpenStudy (anonymous):

@genius12 and the x here is x + 2

OpenStudy (anonymous):

i didn't finished the explanation above -3x=24, x=-24/3=-8